OpenStudy (anonymous):

For the following system, if you isolated y in the first equation to use the Substitution Method, what expression would you substitute into the second equation? 2x + y = 8 −x − y = −5

OpenStudy (experimentx):

y = 8 -2x ---1 -x-y = -5 ---2 -x-(8-2x) = -5 --- solve for x

OpenStudy (anonymous):

oops i forgot the choices! here u go! 8 + 2x 8 – 2x –8 − 2x –8 + 2x

OpenStudy (experimentx):

8 - 2x